What is the standard enthalpy of formation (AHf) in kJ of NO2(g) given: 2 NO28) →N2(g) + 202(g) AHrxn = -67.6 kJ
Answer #1

2 NO2 (g) = N2(g) + 2 O2(g). ∆H = -67.6 KJ

∆H(O2) = 0

∆H(N2) = 0

∆H(reaction) = ∆H(N2) + 2 * ∆H(O2) - 2* ∆H(NO2)

= -67.6 = 0 + 2*0 - 2*(NO2)

NO2 = 33.8 KJ / mol

If you have any questions please comment

If you satisfied with the solution please rate it thanks